# Methods for differential network estimation: an empirical comparison

### v1

@@ 1 Introduction
Within graphical models, we focus specifically on undirected ones. Firstly, directed graphical models are more challenging to estimate (Zheng et al.,, 2018) and additional constraints are required in order to find a unique solution (Giraud,, 2015). Secondly, in many cases, it is of interest to focus only on the presence of a connection, not necessarily its direction.
In the class of undirected graphical models, we consider those associated with the Gaussian distribution. As in many other situations, this particular distribution has many unique theoretical properties that make its use beneficial. In the case of graphical models, it is the fact that Gaussian graph structure can be estimated through a precision (inverse covariance) matrix.
@@
The aim of the article is twofold. First, we will perform an evaluation and a comparison of available methods for differential network estimation through a simulation study. Our focus is on structure learning, so we measure performance by power and false discovery rate related to edge detection. Next, we will investigate the influence of graph structure on methods performance through the same simulation study.
We will start with the various definitions of a differential network. We will then demonstrate the variability of the results of the existing methods for the differential network estimation with a real data example, thus emphasizing the importance of error control. Next, we will describe the estimation methods and compare them in the simulation study, highlighting their advantages and downsides. We will discuss how different graph structures influence the performance of those methods as a class and each of them individually. In the end, we will mention possible directions for future work.
@@ 1.1 What is a differential network?
First, we would like to highlight that multiple definitions of differential networks exist. Let us first define a Gaussian graphical model. We start with a $p$-variate zero mean Gaussian distribution of variables $\boldsymbol{X}=(X_{1},\dots,X_{p})^{T}$ with a covariance matrix $\mathbf{\Sigma}$. Note that we assume that the mean vector is zero without loss of generality, since the covariance structure, modeled by the graphical model, is structurally independent from the mean vector. We can represent this distribution $\mathcal{N}(0,\mathbf{\Sigma})$ in terms of the graph $G=(V,E)$, where $V=\{1,\dots,p\}$ is the set of vertices and $E\subseteq V\times V$ is the set of edges. Each vertex $i$ corresponds to a random variable $X_{i}$, and the absence of an edge between vertices $i$ and $j$ implies conditional independence of the corresponding variables, i.e., $(i,j)\notin EcX_{i}\perp\!\!\!\perp X_{j}|\boldsymbol{X}_{\{l,\,l\neq i,\,j\}}$, where ${\boldsymbol{X}}_{U}$ is a subvector of $\boldsymbol{X}$ for any $U\subset V$. In the Gaussian setting, the support of the precision matrix $\mathbf{\Omega}=\mathbf{\Sigma}^{-1}$ encodes the graph structure. Let $\mathbf{A}$ denote the adjacency matrix of graph G, then $A_{ij}=0$ if and only if $\Omega_{ij}=0$.
Now let us consider two distributions instead of one, $\mathcal{N}(0,\mathbf{\Sigma}^{(1)})$ and $\mathcal{N}(0,\mathbf{\Sigma}^{(2)})$. They can describe two conditions, such as subtypes of a disease or data collected on different equipment or in different laboratories. In this case, it is of interest to investigate the differences between the two distributions. We can do so in terms of graphs again, as supports of precision matrices $\mathbf{\Omega}^{(1)}$ and $\mathbf{\Omega}^{(2)}$ encode graphs’ $G^{(1)}$ and $G^{(2)}$ adjacency matrices $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$. However, how do we define a differential network?
@@
Figure 1 illustrates the three definitions of a differential network. The choice of which definition to use in practical applications should be guided by subject matter considerations. Most methods reviewed here use Definition 1, with the exception of differential connectivity analysis (DCA), that uses Definition 3. Our simulation studies are designed so that differential networks according to the three definitions coincide.
1.2 TCGA breast cancer dataset analysis
Figure 2: Figure demonstrates the variability of results, produced with different methods. Multiplicity of an edge denotes how many methods have identified it and the color corresponds to a method.
Before describing the various methods, we would like to present a real data example to demonstrate how vastly the results vary. In our opinion, this variability highlights the challenge in differential network estimation and emphasizes the need for reliable methods with error control.
In (Koboldt et al.,, 2012), the authors collected 526 samples of patients with breast cancer of four molecular subtypes: luminal A, luminal B, basal-like, and HER2-enriched cancers. The original microarray gene expression dataset has 17 327 genes. For this analysis, we use a pre-processed sub-dataset available in the DiffNetFDR R package (Zhang et al.,, 2019). It consists of two datasets, luminal A subtype with $n_{1}=231$ samples and basal-like subtype with $n_{2}=95$ samples, and $p=139$ genes from the hsa05224 KEGG (Kanehisa and Goto,, 2000) breast cancer pathway.
To identify gene connections that differ between luminal A and basal-like subtypes of breast cancer, we used five different differential network estimation methods. We will describe each method in detail in the following section. Here, however, we would like to focus on the variability of the methods’ results.
To assess the results of the methods based on the different approaches and with various tuning parameters, we have decided to choose the network size as the common denominator. Although it is not a universal practice, aiming for a small network size in the estimation is not uncommon to improve interpretability. Therefore, in this example, we compare five 10-edge networks produced by five different methods. The parameters we used to obtain these estimates are provided in Appendix Appendix A: Real data settings.
Out of the original 139 vertices (genes), 96 are estimated to be isolated by all the methods. The size of the union of all five 10-edge networks is 33 edges. Interestingly, only one edge was identified by all five methods, and only two edges were identified by four methods. Most of the edges were estimated by one method only (see Table 1 and Figure 2). This indicates the difficulty of differential network estimation, and we use this example to highlight it and to motivate our further investigation.
Table 1: A frequency table showing the number of methods, $m$, whose estimated differential networks have $l$ common edges, i.e. no edge was present in all five estimated differential networks, three edges were present in four networks, one edge was estimated by three methods…
Number of methods $m$ 5 4 3 2 1
Number of edges $l$ 1 2 1 6 23
The only edge that was estimated by all five methods is the EGFR-KIT edge. The authors of (Nalwoga et al.,, 2008) and (Kanapathy Pillai et al.,, 2012) report that expression of these genes is associated with basal-like breast cancer subtype which may suggest that this edge is a true differential edge (i.e. present in basal-like type but not in luminal A). The two edges that were estimated by 4 out of 5 methods are IGF1-KIT and EGFR-CDK6. We were not able to find evidence in the literature that these connections may be differential between luminal A and basal-like breast cancer subtypes.
@@ 2 Methods
These methods obtain an estimate of a differential network by testing equality of the entries of a partial correlation matrix (Liu,, 2017), a precision matrix (Xia et al.,, 2015), or an adjacency matrix (Zhao and Shojaie,, 2022) in two conditions.
@@
In our simulation study, we consider representatives of each category. Our choice of specific methods was mainly guided by the software availability: we included only methods that have available implementation in R. An overview of the considered methods is given in Table 2 with a brief description of each.
@@
Name Reference Type Estimand Error control R package Tun. param.
@@
D-trace (Yuan et al.,, 2017) Direct estim. $\mathbf{\Omega}^{(1)}-\mathbf{\Omega}^{(2)}$ No DiffGraph $\lambda$
@@ 2.1 Fused graphical lasso (FGL)
Fused graphical lasso (Danaher et al.,, 2014) is a likelihood-based method for joint estimation of multiple graphical models from datasets with observations belonging to distinct classes. Although it is not the first method in this class – an earlier example can be (Guo et al.,, 2011) – it is currently one of the most mentioned in the literature. The aim is to estimate two graphical models under the assumption that both networks are sparse and at the same time enforcing some similarity between the networks. In the high dimensional case of $p$ bigger than $n_{k}$, $k=1,2$, the authors suggest to maximize the following penalized log-likelihood function:
@@
2.2 Testing equality of partial correlations
The authors of (Liu,, 2017) define a differential network through partial correlations. A partial correlation coefficient of $X_{i}$ and $X_{j}$ given $X_{l\neq i,j}$ can be expressed through precision matrix entries as $\rho_{ij\cdot}=-\Omega_{ij}/\sqrt{\Omega_{ii}\Omega_{jj}}.$ For two conditions, their differential graph is defined similarly to Definition 1 as
@@ 2.2 Testing equality of partial correlations
2.3 Testing equality of the entries of precision matrices
In (Xia et al.,, 2015), a differential network is defined according to Definition 1, as the difference between entries of precision matrices. The authors first perform a test of a global null hypothesis $H_{0}:\mathbf{\Omega}^{(1)}=\mathbf{\Omega}^{(2)}$ and if it is rejected, investigate the structure of the differential network with a multiple testing procedure for a collection of null hypotheses $\left\{H_{ij}:\Omega_{ij}^{(1)}=\Omega_{ij}^{(2)},1\leq i<j\leq p\right\}$. To perform the test of $H_{ij}$, the authors estimate $\mathbf{\Omega}^{(k)}$ by its relation with the coefficients of a set of node-wise linear regression models for $\mathbf{X}^{(k)}$. The test statistics are then obtained as covariances between the residuals of the fitted models.
@@ 2.4 Differential connectivity analysis (DCA)
In our simulation study, we used the authors’ R package DCA available on Github (not to be confused with the package of the same name on CRAN).
@@ 2.5 Direct estimation with lasso penalized D-trace loss (Dtrace)
Similarly to the fused graphical lasso, the authors suggest to use alternating direction method of multipliers to solve the problem.
@@ 3.1 Study set-up
In this section, we introduce our simulation study set-up. To assess the methods’ performance, we decided to vary the structure of the differential network, the sparsity of the graphs, and the number of samples in the data. We elaborate on parameter configurations below.
@@ 3.1.1 Graph generation
All graphs have $p=200$ vertices. The number of edges (size) of $G^{(1)}$ is determined by the number of edges added at each step of the Barabasi-Albert algorithm for scale-free graph generation, one or two. The size of $G^{diff}$ and, consequently, $G^{(2)}$ is determined either by the rewiring procedure or the edge removal procedure. Since the graph generation steps are non-deterministic, they were re-run to achieve the best size consistency across different graph types and sparsity set-ups. The resulting graph sizes are given in Table 3. For simplicity, in the rest of the paper we will refer to the rounded-up graph sizes (50-200, 100-200, 50-400 and 100-400, respectively). We describe the graph generation procedure in detail in Appendix Appendix B: Simulation study data generation. See examples of generated graphs for the 50-200 edge setting in Figure 3 and all graphs in Figure (a) of Supplementary materials.
@@ 3.1.3 Performance evaluation
For each method we vary either the regularization parameter $\lambda$ (or $\lambda_{1}$ and $\lambda_{2}$) or the FDR level $\alpha$ to vary the size of $\hat{E}^{diff}$, from almost complete to almost empty. For each value of the parameter, we compare the estimated differential graph with the true differential graph and calculate the number of true and false positives as
$\begin{split}TP&=|\{(i,j):(i,j)\in E^{diff}\cap\hat{E}^{diff}\}|,\\ FP&=|\{(i,j):(i,j)\in\left(E^{diff}\right)^{c}\cap(i,j)\in\hat{E}^{diff}\}|.\end{split}$
@@
$\begin{split}power&=TP/|E^{diff}|,\\ FDR&=FP/\max(1,FP+TP).\end{split}$
@@
Values of the parameters used in the simulation study are reported in Appendix Appendix C: Simulation settings. Simulation study code is available at github.com/annaplaksienko/Diff_networks_review_simulation.
@@ 3.2 Results
For most settings, all methods exhibit the same trends depending on the graph structure and sparsity. Comparing various differential graph structures (see panels (a), (b) and (c) of Figure 4), we observe that the performance is the best for the random differential graph. It then gets worse for the scale-free graph and finally, the power is very low for almost all methods in all settings for the star differential graph. In other words, the more ”structure” there is in a differential graph, the worse the performance becomes. Moreover, we notice that in most cases, the size of $G^{diff}$ has little effect on performance while an increase in size of $G^{(1)}$ from $\approx 200$ to $\approx 400$ edges influences performance dramatically (compare sets of four facets in each panel and notice how performance drops from the first to the second row). For the random differential graph, the main drop in performance seems to happen when both the size of $G^{(1)}$ and the size of $G^{diff}$ are increased simultaneously. We address these points in the Discussion section.
@@
Figure 4: Estimated power vs FDR for the five considered methods in various sparsity and differential graph structure settings, with varying sample size. Results are averaged over 50 realizations of data.
@@
The best performance overall across all settings is demonstrated by direct estimation with lasso penalized D-trace loss (green ”Dtrace” line in plots). It relies on direct estimation and focuses on the most distinct minimization problem. In most settings the difference in performance may not be large (although still noticeable), but in the 50-200 and 100-200 star differential graph setting it demonstrates a clear advantage compared to other methods.
@@
All testing methods demonstrate very similar performance, with DCA (pink line) providing a slightly higherr power. For very low FDR, all three methods often demonstrate better power than the above-mentioned Dtrace method.
@@
The joint estimation method FGL demonstrates the worst performance. In particular, for the star differential graph, this method has virtually no power at all.

### v2

@@ 1 Introduction
Within graphical models, we focus specifically on undirected ones. Firstly, directed graphical models are more challenging to estimate (Zheng et al.,, 2018) and additional constraints are required to ensure identifiability (Giraud,, 2015). Secondly, in many cases, it is of interest to focus only on the presence of a connection, not necessarily its direction.
In the class of undirected graphical models, we consider those associated with the Gaussian distribution. As in many other situations, this particular distribution has many unique theoretical properties that make its use beneficial. In the case of graphical models, it is the fact that the Gaussian graph structure can be estimated through a precision (inverse covariance) matrix.
@@
The aim of the article is twofold. First, we will perform an evaluation and a comparison of available methods for differential network estimation through a simulation study. Our focus is on structure learning, so we measure performance by power and false discovery rate (FDR) related to edge detection. Next, we will investigate the influence of graph structure on methods performance through the same simulation study.
We will start with the various definitions of a differential network. Next, we will describe the estimation methods and compare them in the simulation study, highlighting their advantages and downsides. We will discuss how different graph structures influence the performance of the methods as a class and each of them individually. Furthermore, we will demonstrate the performance of the six methods in an applied data example, highlighting the variability of result, and in the end, we will discuss our findings and point to possible directions for future work.
@@ 1.1 What is a differential network?
First, we would like to highlight that multiple definitions of differential networks exist. Let us first define a Gaussian graphical model. We start with a $p$-variate zero mean Gaussian distribution of variables $\boldsymbol{X}=(X_{1},\dots,X_{p})^{T}$ with a covariance matrix $\mathbf{\Sigma}$. Note that we assume that the mean vector is zero without loss of generality, since the covariance structure, modeled by the graphical model, is structurally independent from the mean vector. We can represent this distribution $\mathcal{N}(0,\mathbf{\Sigma})$ in terms of the graph $G=(V,E)$, where $V=\{1,\dots,p\}$ is the set of vertices and $E\subseteq V\times V$ is the set of edges. Each vertex $i$ corresponds to a random variable $X_{i}$, and the absence of an edge between vertices $i$ and $j$ implies conditional independence of the corresponding variables, i.e., $(i,j)\notin E$ implies that $X_{i}\perp\!\!\!\perp X_{j}|\boldsymbol{X}_{V\setminus\left\{i,j\right\}}$, where ${\boldsymbol{X}}_{U}$ is a subvector of $\boldsymbol{X}$ for any $U\subset V$. In the Gaussian setting, the support of the precision matrix $\mathbf{\Omega}=\mathbf{\Sigma}^{-1}$ encodes the graphical structure. Let $\mathbf{A}$ denote the adjacency matrix of graph G, then $A_{ij}=0$ if and only if $\Omega_{ij}=0$.
Now let us consider two distributions instead of one, $\mathcal{N}(0,\mathbf{\Sigma}^{(1)})$ and $\mathcal{N}(0,\mathbf{\Sigma}^{(2)})$. They can describe two conditions, such as subtypes of a disease or data collected on different equipment or in different laboratories. In this case, it is of interest to investigate the differences between the two distributions. We can do so in terms of graphs again, as supports of precision matrices $\mathbf{\Omega}^{(1)}$ and $\mathbf{\Omega}^{(2)}$ encode graphs’ $G^{(1)}$ and $G^{(2)}$ and adjacency matrices $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$. However, how do we define a differential network?
@@
Figure 1 illustrates the three definitions of a differential network. The choice of which definition to use in practical applications should be guided by subject matter considerations. Most methods reviewed here use Definition 1, with the exception of differential connectivity analysis (DCA), that uses Definition 3. Our simulation studies are designed so that differential networks according to the three definitions coincide. In this way, each method is used to estimate the type of difference it was originally intended for.
@@ 2 Methods
These methods obtain an estimate of a differential network by testing equality of the entries of a partial correlation matrix (Liu,, 2017), a precision matrix (Xia et al.,, 2015; Chen et al.,, 2024), or an adjacency matrix (Zhao and Shojaie,, 2022) under two conditions.
@@
In our simulation study, we consider representatives of each category. Our choice of specific methods was mainly guided by the software availability: we included only methods that have an available implementation in R. An overview of the considered methods is given in Table 1 with a brief description of each.
@@
Name Ref. Type Estimand Error control R package Tun. param.
@@
ONDSA (Chen et al.,, 2024) Testing $E^{diff}=\left\{(i,j):\Omega_{ij}^{(1)}\neq\Omega_{ij}^{(2)}\right\}$ Asymp. FDR ONDSA No
@@
DTrace (Yuan et al.,, 2017) Direct estim. $\mathbf{\Omega}^{(1)}-\mathbf{\Omega}^{(2)}$ No DiffGraph $\lambda$
@@ 2.1 Fused graphical lasso (FGL)
The fused graphical lasso (Danaher et al.,, 2014) is a likelihood-based method for the joint estimation of multiple graphical models from datasets with observations belonging to distinct classes. Although it is not the first method in this class – an earlier example can be (Guo et al.,, 2011) – it is currently one of the most mentioned in the literature. The aim is to estimate two graphical models under the assumption that both networks are sparse, while at the same time enforcing some similarity between the networks. In the high-dimensional case of $p$ bigger than $n_{k}$, $k=1,2$, the authors suggest to maximize the following penalized log-likelihood function:
@@
2.2 Testing the equality of partial correlations (PCor)
The authors of (Liu,, 2017) defines a differential network through partial correlations. A partial correlation coefficient of $X_{i}$ and $X_{j}$ given $X_{l\neq i,j}$ can be expressed through precision matrix entries as $\rho_{ij\cdot}=-\Omega_{ij}/\sqrt{\Omega_{ii}\Omega_{jj}}.$ For two conditions, their differential graph is defined similarly to Definition 1 as
@@ 2.2 Testing the equality of partial correlations (PCor)
2.3 Testing the equality of precision matrices (PMat)
In (Xia et al.,, 2015), a differential network is defined according to Definition 1, as the difference between entries of precision matrices. The authors first perform a test of a global null hypothesis $H_{0}:\mathbf{\Omega}^{(1)}=\mathbf{\Omega}^{(2)}$ and if it is rejected, investigate the structure of the differential network with a multiple testing procedure for a collection of null hypotheses $\left\{H_{ij}:\Omega_{ij}^{(1)}=\Omega_{ij}^{(2)},1\leq i<j\leq p\right\}$. To perform the test of $H_{ij}$, the authors estimate $\mathbf{\Omega}^{(k)}$ by its relation with the coefficients of a set of node-wise linear regression models for $\mathbf{X}^{(k)}$. The test statistics are then obtained based on a bias-corrected estimator of the covariances between the residuals of the fitted models.
@@ 2.3 Testing the equality of precision matrices (PMat)
2.4 Omics Networks Differential and Similarity Analysis (ONDSA)
In (Chen et al.,, 2024), a differential network is also defined by the differences in precision matrices (Definition 1) and differential and similar structures are identified in a two-step procedure with FDR control. First, individual networks are estimated via (Ren et al.,, 2015). In order to estimate individual entries $\mathbf{\Omega}_{ij}$, ($X_{i}$, $X_{j}$) is regressed on the remaining variables by the scaled lasso. The scaled lasso is ”tuning-free”, as the authors suggest to fix $\lambda=\sqrt{(2/N)\log(p/\sqrt{N})}$. The residuals from this regression is then used to estimate $\mathbf{\Omega}_{ij}$. Then, the null hypothesis $H_{0}:\mathbf{\Omega}^{(1)}_{ij}=\mathbf{\Omega}^{(2)}_{ij}$ is tested for each edge to construct the set of differential edges $E^{diff}$. To correct for correlations among entry-wise test statistics caused by overlapping nodes in the network, a conditional FDR adjustment is applied, following the recommendation of (Efron,, 2007). Furthermore, on the set complementary to the estimated differential edges $({E^{diff}})^{c}$, this method applies another testing procedure to divide the remaining edges into common nonzero edges and common zero edges (absence of edges).
In our simulations, we used the authors’ package ONDSA, available on Github.
@@ 2.5 Differential connectivity analysis (DCA)
In our simulation study, we used the authors’ R package DCA available on Github (not to be confused with the package of the same name on CRAN). We have added correction for multiplicity with the Benjamini-Hochberg procedure.
@@ 2.6 Direct estimation with lasso penalized D-trace loss (DTrace)
Similarly to the fused graphical lasso, the authors suggest to use alternating direction method of multipliers to solve the problem and obtain $\hat{\mathbf{\Delta}}$.
@@ 3.1 Study set-up
In this section, we introduce our simulation study set-up. To assess the methods’ performance, we decided to vary the structure of the differential network, the sparsity of the graphs, and the sample size. We elaborate on parameter configurations below.
@@ 3.1.1 Graph generation
All graphs have $p=200$ vertices. The number of edges (size) of $G^{(1)}$ is determined by the number of edges added at each step of the Barabasi-Albert algorithm for scale-free graph generation, one or two. The size of $G^{diff}$ and, consequently, $G^{(2)}$ is determined either by the rewiring procedure or the edge removal procedure. Since the graph generation steps are non-deterministic, they were re-run to achieve the best size consistency across different graph types and sparsity set-ups. The resulting graph sizes are given in Table 2. For simplicity, in the rest of the paper we will refer to the rounded-up graph sizes (50-200, 100-200, 50-400 and 100-400, respectively). We describe the graph generation procedure in detail in Appendix B. See examples of generated graphs for the 50-200 edge setting in Figure 2 and all graphs in Figure (a) of Supplementary materials.
@@ 3.1.3 Performance evaluation
For each method we vary either the regularization parameter $\lambda$ (or $\lambda_{1}$ and $\lambda_{2}$) or the FDR level $\alpha$ to vary the size of $\hat{E}^{diff}$, from almost complete to almost empty. For each value of the parameter, we compare the estimated differential graph with the true differential graph and calculate the number of true and false positive edges as
$\begin{split}TP&=|\{(i,j):(i,j)\in E^{diff}\cap\hat{E}^{diff}\}|,\\ FP&=|\{(i,j):(i,j)\in\left(E^{diff}\right)^{c}\cap\hat{E}^{diff}\}|.\end{split}$
@@
$\begin{split}\mbox{power}&=TP/|E^{diff}|,\\ \mbox{FDR}&=FP/\max(1,FP+TP).\end{split}$
@@
Values of the parameters used in the simulation study are reported in Appendix C. Simulation study code is available at github.com/annaplaksienko/Diff_networks_review_simulation.
@@ 3.2 Results
Figure 3: Estimated power vs FDR for the six considered methods in various sparsity and differential graph structure settings, with varying sample size. Results are averaged over 50 realizations of data.
@@
For most settings, all methods exhibit the same trends depending on the graph structure and sparsity. Comparing various differential graph structures (see panels (a), (b) and (c) of Figure 3), we observe that overall, performance is similar for the random and the scale-free graphs. However, for the star differential graph the power is very low for almost all methods in all settings. In other words, the more ”structure” there is in a differential graph, the worse the performance becomes. Moreover, we notice that in most cases, the size of $G^{diff}$ has little effect on performance (compare left and right facets in each panel) while an increase in size of $G^{(1)}$ from $\approx 200$ to $\approx 400$ edges influences performance dramatically (compare the first and the second row in each panel and notice the drop in performance).
@@
The best overall performance is demonstrated by DTrace (green line in the plots). It relies on direct estimation and focuses on the most distinct minimization problem. In most settings the difference in performance may not be large (although still noticeable), but in the 50-200 and 100-200 star differential graph setting it demonstrates a clear advantage compared to most other methods.
Testing methods PMat (blue line) and PCor (turquoise line) demonstrate very similar performance under most settings, with an exception of the scale-free differential graph, 50-200 and 100-200 edges with 100 samples, where PCor demonstrates noticeably superior performance. DCA (magenta line) usually is either on par with or slightly outperforms both. ONDSA (yellow line) is inconsistent, showing in some cases inferior performance to all above mentioned testing methods, and in some cases noticeably outperforming all of them – see scale-free differential graph, 50-200 and 100-200 edges with 100 samples and especially star differential graph, 50-200 and 100-200 edges with 400 samples.
@@
For very low FDR, all testing methods often demonstrate better power than the above-mentioned DTrace method.
@@
The joint estimation method FGL demonstrates the worst performance. In particular, for the star differential graph, this method has no power at all.
4 Real data application
Before moving on to the Discussion, we would like to also compare the presented methods on real data. This example, on the one hand, highlights the variability of results produced by different methods and hence emphasizes the difficulty of the problem, and on the other hand, provides evidence that the presented methods are capable of detecting known biological phenomena and should, therefore, be further developed and perfected.
Figure 4: The figure demonstrates the variability of results produced with different methods. The multiplicity of an edge denotes how many methods have identified it and the color corresponds to a method.
We consider a dataset from (Koboldt et al.,, 2012), where the authors collected 526 samples of patients with breast cancer of four molecular subtypes: luminal A, luminal B, basal-like, and HER2-enriched cancers. The original microarray gene expression dataset has 17 327 genes. For this analysis, we use a pre-processed sub-dataset available in the DiffNetFDR R package (Zhang et al.,, 2019). It consists of data from two of the subtypes, the luminal A subtype with $n_{1}=231$ samples and the basal-like subtype with $n_{2}=95$ samples, and $p=139$ genes from the hsa05224 KEGG (Kanehisa and Goto,, 2000) breast cancer pathway.
To identify gene connections that differ between luminal A and basal-like subtypes of breast cancer, we used all the six differential network estimation methods that we discussed above. For the testing methods, we used $\alpha=0.05$, while for FGL and DTrace, the tuning parameters were estimated with the Bayesian Information Criterion (BIC). All values are provided in Appendix A. The different methods produced networks of the following sizes (in decreasing order): PMat – 18 edges, DCA and FGL – 10 edges, PCor – 8 edges, DTrace – 5 edges, and ONDSA – 4 edges.
To compare the results, we constructed a union of all six estimated networks. Out of the original 139 vertices (genes), 92 are estimated to be isolated by all the methods. The size of the union of all six networks is 36 edges. Interestingly, only one edge was identified by all six methods (EGFR-KIT), by five out of six (KIT-IGF1) and by four out of six (Wnt10a-PIK3CD), respectively. Most of the edges were estimated by one method only (see Table 3 and Figure 4).
Table 3: A frequency table showing the number of methods, $m$, whose estimated differential networks have $l$ common edges, i.e. one edge was present in all six estimated differential networks, one edge was present in five networks, one edge was estimated by four methods …
Number of methods $m$ 6 5 4 3 2 1
Number of edges $l$ 1 1 1 1 5 27
The only edge estimated by all six methods is the EGFR-KIT edge. The authors of (Nalwoga et al.,, 2008) and (Kanapathy Pillai et al.,, 2012) report that expression of both of these genes is associated with the basal-like breast cancer subtype which suggests that this edge is a true differential edge, i.e. present in the basal-like subtype but not in luminal A. If we consider each node individually, (Kashiwagi et al.,, 2013) identifies c-KIT as a potential biomarker and possible molecular target for therapy for patients with basal-like breast cancer. For EGFR, also associated to the basal-like subtype, there has been arguments regarding if it is a potential therapeutic target (Siziopikou and Cobleigh,, 2007) or ”just” a biomarker (Burness et al.,, 2010).
We were not able to find specific literature support for the KIT-IGF1 edge, the edge estimated by five out of six methods. However, we already mentioned above that c-KIT has been associated with basal-like breast cancer and similar evidence can be found for IGF1. The authors of (Christopoulos et al.,, 2015) mention that basal-like tumors frequently induce IGF1 transcriptional upregulation and studies (Klinakis et al.,, 2009; Kumar and Chaudhri,, 2024) identify IGF1 as a therapeutic target specifically in basal-like breast cancer.
The edge Wnt10a-PIK3CD was estimated by four out of six methods. In (Li et al.,, 2024) it was discovered that Wnt10a is strongly linked to the basal-like subtype, while the authors of (He et al.,, 2023) hypothesized that PIK3CD is associated with a favorable prognostic biomarker for overall survival in the luminal A subtype.
We consider it promising that edges with known biological support were identified by all/most methods, even if some evidence is circumstantial. Regarding the ”single” edges, i.e. those identified only by one of the methods, it is unclear whether those are false positives or if different methods capture different parts of the true graph.